A particle moves under a central force field. The quantity  [d2/dθ2 (1/r)]  is proportional to 1/r. The symbols have standard meaning. The magnitude of the force is proportional to


(a) 1/r2


(b) 1/r


(c) 1/ r3


(d) r


(e) none of these


1
Expert's answer
2021-03-09T13:18:39-0500

Given,

"\frac{d^2}{d\theta^2}(\frac{1}{r})=\frac{1}{r}"


Let "\frac{1}{r}=k"


"\frac{d^2k}{d\theta^2}=k"


we can write it as "\frac{d^2k }{k}= d\theta^2"

Now, "\frac{d^2k }{kdt^2}= \frac{d\theta^2}{dt^2}"

Angular acceleration "(\alpha)\propto r\frac{d^2(1/r)}{dt^2}"

Hence, we can conclude that it is proportional to r hence correct answer be option (d)